A ball is thrown upward from the top of a building 80 feet tall with an initial velocity of 64 feet per second. Solve the equation: -16t^2+64t+80=144 to find the time t​ (in seconds) at which the ball is 144 feet from the ground.
Answer by ankor@dixie-net.com(22740)

A ball is thrown upward from the top of a building 80 feet tall with an initial velocity of 64 feet per second.
Solve the equation: -16t^2+64t+80=144 to find the time t​ (in seconds) at which the ball is 144 feet from the ground.
:
-16t^2 + 64t + 80 = 144
Simplify, divide equation by -16
t^2 - 4t - 5 = -9
t^2 - 4t - 5 + 9 = 0
t^2 - 4t + 4 = 0
Factors to
(t-2)(t-2) = 0
t = 2 Seconds